1)an arithmetic progression has a first term of 32 , a 5th term of 22 and a last term of -28 .find the sum of all the terms in the progression .
2)each year a school allocates a sum of money for the library . the amount allocated each year increases by 2.5 % of the amount allocated the previous year . in 2005 the school allocated $2000.find the total amount allocated in the years 2005 to 2014 inclusive
1
Expert's answer
2020-05-26T20:15:54-0400

1.)

The first term in arithmetic progression is "(a) = 32"

Since the "5^{th}" term in the arithmetic progression is "22" so we can write:

"t_5 = 22 \\
\Rightarrow a + (5-1)d = 22" "\hspace{1 cm} [ \because" "t_n = a + (n-1)d \, \," ( Where a = first term and d = common difference) ]

"\Rightarrow a + 4d = 22 \\
\Rightarrow 32 + 4d = 22 \\
\Rightarrow 4d = 22 - 32 \\"

"\Rightarrow d = - \frac{10}{4} \\
\Rightarrow d = - 2.5"

Let us assume that number of terms in this progression be "n"

Now:

Since the last term of this progression is "-28" so we can write:

"t_n = -28 \\
\Rightarrow a + (n-1)d = -28 \\
\Rightarrow 32 + (n-1)(-2.5) = -28 \hspace{1 cm}" [ "\because a = 32 \, \,\, and \,\,\, d = -2.5" ]

"\Rightarrow n-1 = \frac{ -28 - 32}{-2.5} \\
\Rightarrow n - 1 = 24 \\
\Rightarrow n = 25"

Hence:

The sum of all terms in this progression is:

"S_n = \frac{n}{2} (a + l) \hspace{1 cm}" [ Where a = first term and l = last term ]

"\hspace{0.5 cm} = \frac{25}{2} (32 -28) \\
\hspace {0.5 cm} = 50"

2.)

Since the amount of money each year increased by "2.5 \%" of the amount of money allocated in the previous year so each year added a fixed amount money of the money allocated in the previous year and allocated money from "2005 \, \text{to} \, 2014" forms an arithmetic progression.

Now:

Amount of money allocated in "2005" is "\$ \, 2000" which is also a first term (a) in this progression.

Number of terms in this progression is:

"n =" number of years from "2005 \, \text{to} \, 2014 = 10"

Amount of money allocated in 2006 is:

"= \$ \, \left( 2000 + 2000 \times 2.5 \% \right) \\
= \$ \, 2050"

So, common difference in this progression is:

"d = 2050 - 2000 = 50" [ "\because" Common difference in an arithmetic progression (d) = next term - previous term ]

Hence:

The total amount of money allocated from "2005 \, \text{to} \, 2014" inclusive is:

"= \frac{n}{2} \left( 2a + (n-1)d \right) \\"

"= \frac{10}{2} \left( 2(2000) + (10-1)50 \right) \\
= \$ \, 22250"